OpenStudy (anonymous):

The chess club has 18 students. How many different teams of 4 can be formed for competition? 3,060 18,360 73,440

OpenStudy (anonymous):

would it be 73440

jimthompson5910 (jim_thompson5910):

That would be the correct answer if order mattered. However, I don't think order matters.

OpenStudy (anonymous):

then wht do u think

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

Use the combination formula n C r = (n!)/(r!(n-r)!) instead of the permutation formula

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 18 C 4 = (18!)/(4!*(18-4)!) 18 C 4 = (18!)/(4!*14!) 18 C 4 = (18*17*16*15*14!)/(4!*14!) 18 C 4 = (18*17*16*15)/(4!) 18 C 4 = (18*17*16*15)/(4*3*2*1) 18 C 4 = (73440)/(24) 18 C 4 = 3060 So there are 3060 different teams